20 bags of cement, of 40 kg each, cost Rs1800. Find the cost of 15 bags of cement weighing 3 kg each.

Grade:6

2 Answers

Rajat Meshram
6 years ago
20 bags ,each 40 kg total-800 kg1 kg cost = 1800/800=2.25rs15 bags 3 kg -45 kgstotal cost -45 * 2.25= 101.25
Om
5 years ago
Dear student 
20 bags of 40 kg means a total of 800 kg costs Rs1800
So 1kg cost is Rs 2.25
15 bags of 3 kg means a total of 45 kg which will cost Rs(2.25x45) = Rs101.25
I hope this answer helps you
With regards 
A fellow student